Please help me solve the following problem:
State the x-intercepts of the graph of the equation. Then find the coordinates of the vertex.
y=(x+3)(x-5)
: Please help me solve the following problem:
State the x-intercepts of the graph of the equation. Then find the coordinates of the vertex.
y=(x+3)(x-5)

Answer by scott8148(2759)

when the graph crosses the x-axis (x-intercept), the value of y is zero
__ this happens when x+3=0 or x-5=0

so the x-intercepts are -3 and 5

the x-coordinate of the vertex is midway between the intercepts (on the axis of symmetry)
__ midway between -3 and 5 is 1 ([5-3]/2)

substituting 1 for x in the equation gives y=(1+3)(1-5) or y=-16

so the vertex is at (1,-16)